OpenStudy (anonymous):

Write an expression. Ann is 20 years older than Bill. Carmen is 4/5 as old as Ann. Dana is four years older than Ann. If Bill is B years old, find Carmen's age. Find Dana's Age. Find Ann's age.

OpenStudy (anonymous):

Ann is 20 years older than bill. so if bill was 10 years old, ann would be 30 years old suppose ann was A years old. write a mathematical expression for A

OpenStudy (sriram):

ann=b+20 C=(b+20)*4/5 d=b+24

OpenStudy (anonymous):

Thanks so much. I was close, lol!